Solve 2x^2-16x+20=132 by completing the square.
Answer by mananth(16949)

2x^2-16x+20=132
2x^2-16x=112
/2
x^2-8x=56
x^2-8x+16=56+16
(x-4)^2= 72
take the square root on both sides
(x-4)= +/-sqrt72
x= 4+sqrt72 =12.48
OR
x=4-sqrt72 = -4.48